The graph of the relation x^2 - 4y^2 + 2x - 8y - 10 = 0 is a (an)
a. point
b. straight line
c. pair of intersecting straight lines
d. circle
e. ellipse
f. hyperbola
g. parabola
h. none of these

Answer by algebrapro18(249)

You need to complete the square to get this equation into standard form and then see what it matches up to.

x^2 - 4y^2 + 2x - 8y - 10 = 0

x^2 + 2x - 4(y^2 - 2y) = 10

x^2 + 2x + 1 - 4(y^2 - 2y + 1) = 10+4+1

(x+1)^2 - 4(y-1)^2 = 15

((x+1)^2)/15 - (4(y-1)^2)/15 = 1

I can't remember if this is the standard form for an ellipse or a hyperbola but it is one of the two.